Assume that two identical firms in a purely oligopolistic industry selling a homogenous product agree to share the maket equally. The total market demand function for the commodity is Qd = 240 – 10P. The cost schedules of the firms are given in the following table:

q1 40 50 60 80 q2 50 70 100
SMC1 (Rs.) 8 10 12 16 SMC1 (Rs.) 4 6 9
SAC1 (Rs.) 13 12.3 12 13 SAC1 (Rs.) 7 6 7

Question 1: Profits for this firm will be:
a. Rs. 420
b. Rs. 130 (wrong)
c. Rs. 350
d. Rs. 450
Question 2: When q1 = 40, What will be MR1?
a. 2
b. 8 ✓
c. 5
d. 4
Question 3: When q1 = 40, what will be the profit maximising output for the first firm?
a. 30
b. 60
c. 40 ✓
d. 20
Question 4: When q1 = 50, what will be MR1?
a. 7
b. 2
c. 4 ✓
d. 3
Question 5: When q1 = 60, what will be MR1?
a. 0 ✓
b. 2
c. 4
d. 6
Question 6: When q1 = 80, what will be MR1?
a. 7
b. -4
c. 5
d. -8 ✓
✓ mark are the right answers. Please give explanation for this problem
1
Expert's answer
2020-05-19T10:28:10-0400

Question 1: Profits for this firm will be:

P = 24 - 0.1Q,

q1 = q2 = 50, so Q = 100 units.

P = 24 - 0.1×100 = 14.

TP2 = (14 - 7)×50 = 350.

b. Rs. 130 (wrong)

c. Rs. 350 ✓


Question 2: When q1 = 40, What will be MR1?

MR = TR'(Q) = 24 - 0.2Q,

Q = 40×2 = 80, so:

MR1 = 24 - 0.2×80 = 8.

b. 8 ✓


Question 3: When q1 = 40, what will be the profit maximising output for the first firm?

In this case q1 = q2 = 40, Q = 80.

c. 40 ✓


Question 4: When q1 = 50, what will be MR1?

MR1 = 24 - 0.2×(50×2) = 4.

c. 4 ✓


Question 5: When q1 = 60, what will be MR1?

MR1 = 24 - 0.2×(60×2) = 0

a. 0 ✓


Question 6: When q1 = 80, what will be MR1?

MR1 = 24 - 0.2×(80×2) = -8.

d. -8 ✓